The wire of an electric heater (100w, 200v) is cut from middle into two equal parts. These parts are connected together in parallel and then connected to the same 200 v source. The power liberated is now-

Grade:12

1 Answers

Arun
6 years ago
Dear Deepak

power rating,m P = 100 W
voltage supply, V = 200V
thus resistance of the coil:
P = V2 / R
thus R = 40000/100 = 400 ohm
hence of the two new parts produced, new resistance of each part, R' = 200 ohm each
on parallel connection, net resistance:
1/Rt = 1/R' + 1/R'
hence Rt = R'/2 = 100 ohm
and the new power rating, P = V2 /Rt = 40000/100 = 400 W
 
Regards
Arun (askIITians forum expert)